OpenStudy (anonymous):

tan5x +tan x/1-tan5xtanx

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

just notice \[\tan(a+b)=\frac{\tan a+\tan b}{1-tana \tan b}\]

OpenStudy (anonymous):

a. tan 5x b. tan x c. tan 4x d. tan 6x which one is the answer? to the original problem

OpenStudy (anonymous):

please and thank you

OpenStudy (anonymous):

try with hint let a=5x and b=x

OpenStudy (anonymous):

what will u get?

OpenStudy (anonymous):

d?

OpenStudy (anonymous):

thats right

OpenStudy (anonymous):

Write sin3x cos4x + cos3x sin4x in terms of a single trigonometric function.

OpenStudy (anonymous):

hints please

OpenStudy (anonymous):

\[\sin(a+b)=\sin a \cos b+\cos a \sin b\]

OpenStudy (anonymous):

a. cos 7x b. sin(-x) c. sin 7x d. cos(-x) is the answer c?

OpenStudy (anonymous):

i meant b

OpenStudy (anonymous):

c is right

OpenStudy (anonymous):

simplfy cos [theta + pi/3]?

OpenStudy (anonymous):

\[\cos (a+b)=\cos a \cos b-\sin a \sin b\]

OpenStudy (anonymous):

\[\tan 5\pi/6 +\tan \pi/6 \over 1 - \tan 5\pi/6\tan \pi/6\]

OpenStudy (anonymous):

is it costheta + 1/2? to the simplify answer?

OpenStudy (anonymous):

\[\cos(\theta+\frac{\pi}{3})=\frac{1}{2}\cos \theta-\frac{\sqrt{3}}{2} \sin \theta\]

OpenStudy (anonymous):

given sin 4/5 in quadrant 1 and cos -12/13 in quadrant 2 find cos (a+b)

OpenStudy (anonymous):

a. -16/65 b. -56/65 c. 16/65 d. 56/65

OpenStudy (anonymous):

@mathwhiz99 @lgbasallote @Hero @KingGeorge @satellite73 One of them can help you when they get online, I'm sure :)

OpenStudy (anonymous):

somone hurry help answer this

OpenStudy (anonymous):

I think you meant sin(a) = 4/5 and a is in Q1 cos(b) = -12/13 and b is in Q2 Since a is in Q1, cos(a)>0, and since b is in Q2, sin(b)>0 Therefore, sin(a) = 4/5 => cos(a) = 3/5 and cos(b) = -12/13 => sin(b) = 5/13 Cos(a + b) = cos(a)cos(b) - sin(a)sin(b) = (3/5)(-12/13) - (4/5)(5/13) = (-36/65) - (20/65) = -56/65 The answer is b.